OpenStudy (anonymous):

If r = 9, b = 5, and g = -6, what does (r + b - g)(b + g) equal? -20 -8 8 20

rebeccaxhawaii (rebeccaxhawaii):

plug in your information

Directrix (directrix):

r = 9, b = 5, and g = -6 =============== (r + b - g) * (b + g) = [( 9 + 5 - ( - 6 ) ] * [ 5 + (-6) ] = @WhatTehHeck Do you want to give this a go? Add the numbers within each bracket and then multiply those results together.

Directrix (directrix):

[( 9 + 5 - ( - 6 ) ] = [ 14 + 6] = 20 [ 5 + (-6) ] = [ -1 ] = -1 ------------------------- [( 9 + 5 - ( - 6 ) ] * [ 5 + (-6) ] = 20* (-1) = -20